# Dynamics of “Spontaneous” Topic Changes in Next Token Prediction with Self-Attention

### v2

@@ 5 Explaining Topic Shifts
Definition 4 defines an ambiguous sequence as one where the highest-probability next-token predictions include tokens from both within and outside the input topic, reflecting natural ambiguity from overlapping topics. For instance, the input sequence ${\bf X}=[{\bf e}_{1},{\bf e}_{4},{\bf e}_{1},{\bf e}_{4}]^{\top}$ in Figure 4 (middle) belongs to both $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$. $\widehat{\mathcal{G}}_{a}^{(4)}({\bf X})$ is $\{{\bf e}_{1}\}$, as depicted in $\mathcal{G}_{a}^{(4)}$ from Figure 3 (left) and $\widehat{\mathcal{G}}_{ab}^{(4)}({\bf X})$ is $\{{\bf e}_{1},{\bf e}_{4}\}$, as shown in $\mathcal{G}_{ab}^{(4)}$ from Figure 3 (right). $\widehat{\mathcal{G}}_{a}^{(4)}({\bf X})$ is the subset of $\widehat{\mathcal{G}}_{ab}^{(4)}({\bf X})$, though $\hat{y}_{{\bf w}_{ab}}\not\in\widehat{\mathcal{G}}_{a}^{(4)}({\bf X})$. We can argue that the next token predicted from an ambiguous sequence cannot be considered as a topic change, as it lacks the clear trigger phenomenon observed in human cognition. To address this, we propose a formal definition for a topic change:
@@
In Figure 4, ${\bf W}_{ab}$ for the last input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{4},{\bf e}_{4},{\bf e}_{4}]^{\top}$ can be considered as change of topic.
Building on the formal definitions of topic continuity, ambiguous sequences, and topic changes, we now present a necessary condition for a sequence to induce a topic change. This is achieved by introducing our final definition, grounded in the highest-priority SCC as determined by the order in the attention layer. Now, our second key result.
@@
Under the same settings and assumptions on datasets and training in Theorem 3, let ${\bf X}$ be a sequence that belongs to $\mathbb{T}_{a}$. If $\tilde{\bf W}_{ab}$ changes topic $\mathbb{T}_{a}$ for $\bf X$ then $\exists x_{j}\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$ such that $\forall x_{i}\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$, the number of times ${\bf x}_{j}$ appears in $\bf X$ is greater than the number of times ${\bf x}_{i}$ appears in $\bf X$.
Theorem 4 implies that, for a given sequence X from $\mathbb{T}_{a}$ and its corresponding TPG, a necessary condition for a topic change is the presence of a lower-priority token that appears more frequently than any of the higher-priority tokens. This can be intuitively understood through our analogy: if the oracle is following a conversation on Topic A but the conversation contains repeated components with lower importance in Topic A, its knowledge of Topic B may steer the response toward Topic B, thereby initiating a shift away from Topic A. A natural question arises: what do these findings imply in practice? Specifically, how does the probability of change of topic behaves as the input sequence length increases or as the ambiguity between topics grows? The following theorem provides insights into these dynamics.
@@
    If $p>\max_{x\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})}\mathbb{P}\left({\bf x}_{i}^{\prime}={\bf x}\right)$ then
    $\mathbb{P}(\tilde{\bf W}_{ab}$ changes topic $\mathbb{T}_{a}$ for sequence ${\bf X}^{\prime})\to 0$ as $T\to\infty.$
@@
    If $l$ increases then the probability that $\exists x_{j}^{\prime}\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$ such that $\forall x_{i}^{\prime}\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$, the number of times ${\bf x}_{j}^{\prime}$ appears in $\bf X^{\prime}$ is greater than the number of times ${\bf x}_{i}^{\prime}$ appears in $\bf X^{\prime}$ does not increase.
@@
(a) Length of sequences.
(b) Overlap between topics.
Figure 6: The proportion of keep topic (topic continuity), ambiguity (ambiguous sequence), and change topic (change of topic) as (a) the input sequence length increases and (b) the overlap between topics increases.
@@
To illustrate Theorem 5, we conduct simulations following the same implementation settings as described in Section 4. We approximate $\tilde{\mathbf{W}}_{a}$ and $\tilde{\mathbf{W}}_{ab}$ as the results obtained after $\tau=8000$ iterations of Algo-GD. We quantify the proportion of test sequences in which $\tilde{{\bf W}}_{ab}$ keeps $\mathbb{T}_{a}$ (keep topic), proportion of ambiguous sequences in $\mathbb{T}_{ab}$ (ambiguity) and proportion of times in which $\tilde{{\bf W}}_{ab}$ changes topic (change topic). First, we explore the effect of longer sequences by varying the length $T$ of the test sequences $\bf Z$. We increase $T$ from 4 to 512. Figure 6(a) illustrates how the proportion of change topic decreases as $T$ increases. Second, we investigate the effect of overlapping topics by generating TPGs with an increased number of edges $L$. Intuitively, a higher $L$ results in an increase $l$ and a greater overlap between TPGs of different topics. We vary $L$ from 4 to 18. Figure 6(b) demonstrates that as $L$ increases, ambiguity increases, while proportion of change topic decreases. These two findings contrast with expectations derived from human cognition but align with the result of Theorem 5. Lastly, among the 85000 test sequences generated for these experiments, 99.98% satisfy Theorem 4 (i.e., topic changes occur when a low-priority token appears more frequently than high-priority tokens). The remaining 0.02% mismatched cases were solely due to minor approximation discrepancies in the attention softmax. These results validate Theorem 4 (simulation details in Appendix C).
Related Work
@@
Training and generalization of Transformer. (1) Properties of Self-attention and Softmax. The self-attention mechanism employs the softmax function to selectively emphasize different parts of the input. Gu et al., (2024), Goodfellow et al., (2016), and Deng et al., (2023) underscore the pivotal role of the softmax function in shaping attention distributions, influencing how models process and prioritize information within input sequences. Bombari and Mondelli, (2024) examined the word sensitivity of attention layers, revealing that softmax-based attention-layers are adept at capturing the significance of individual words. Despite these strengths, recent researches have highlighted limitations of the softmax function (Saratchandran et al.,, 2024; Deng et al.,, 2023). (2) Optimization in attention-based models. Additionally, recent researches interpret Transformer models as kernel machines, akin to support vector machines (SVMs), with self-attention layers performing maximum margin separation in the token space (Tarzanagh et al., 2023a, ; Tarzanagh et al., 2023b, ; Li et al., 2024b, ; Julistiono et al.,, 2024). (3) Chain-of-Thought and In-Context Learning. Moreover, transformers exhibit remarkable abilities in generalization through in-context learning (ICL), where models effectively learn from contextual cues during inference (Brown et al.,, 2020; Xie et al.,, 2022; Olsson et al.,, 2022). Chain-of-Thought (CoT) prompting (Wei et al.,, 2022; Zhou et al.,, 2023; Shao et al.,, 2023; Li et al., 2024a, ) enhances this by breaking down reasoning processes into intermediate steps, highlighting the emergent reasoning abilities of transformers. (4) Improvement efficiency of transformers. Recent advancements aim to improve the computational efficiency of transformers (Kitaev et al.,, 2020; Choromanski et al.,, 2021; Sukhbaatar et al.,, 2019; Wang et al.,, 2020), ensuring their viability for large-scale deployment while maintaining or enhancing their representational capabilities.
@@
Self-Attention and topic dynamics. Advancements in the field has significantly advanced our understanding of how self-attention mechanisms in transformers manage evolving semantic contexts, with works have explored various aspects of topic modeling, including dynamic topic structures (Miyamoto et al.,, 2023), hierarchical relationships (Lin et al.,, 2024), topic-aware attention (Panwar et al.,, 2021), and mechanistic analysis of topic structure (Li et al.,, 2023). While these works provide insights into managing static and hierarchical topic structures, our work focuses on the topic changes during next-token prediction within the given input sequences from a specific topic.
@@ 6 Discussion
Our exploration with a simple single-layer attention model reveals fundamental clues on the distinctions between human and AI conversational dynamics. While humans naturally shift topics through curiosity or contextual triggers, the model remains anchored to the context. These observations suggest that human-like spontaneity in topic changes demands mechanisms beyond pattern recognition, raising deeper questions about the nature of interaction and highlighting the importance of creativity in advancing conversational AI.
However, our analysis relies on the simplified single-layer self-attention setting from Li et al., 2024b . This does not directly extend to complex modern LLMs with deeper architectures and varied training goals. Future work will explore how these findings translate to more intricate models, examining their ability to change topics. Moreover, the recurring theme of topic continuity characterized in this work opens new avenues for applications, such as automatic change topic detection in conversational AI based on discrepancies between model predictions and speaker outputs.
@@ Appendix A Technical Proofs
A.1 Proofs of Lemmas
A.1.1 Proof of Lemma 2
@@ Appendix B Detailed Explanation of Assumption 4
Figure 7: Illustration of Assumption 4. Here are two datasets related to the TPGs, $\mathcal{G}_{a}^{(4)}$ and $\mathcal{G}_{b}^{(4)}$, from Figure 3. From the directed arrows in $\mathcal{G}_{a}^{(4)}$, we can generate a dataset with the last token $\bf{e}_{4}$ for $\mathbb{T}_{a}$, which can reconstruct back the $\mathcal{G}_{a}^{(4)}$. A similar process applies for the $\mathcal{G}_{b}^{(4)}$.
Appendix C Detailed Simulation Studies
C.1 Simulation Process
@@
C.2 Experiments Results
In this section, we provide empirical evidence to support our findings outlined in Theorem 3, Theorem 4, and Theorem 5, offering detailed insights into the behavior of self-attention mechanisms for changing topics.
C.2.1 Simulation in Section 4
For Figure 5 we used $K=10,L=4,d=16,\text{and }T_{test}=T_{train}=4$, we also showed point-wise confidence intervals of rank similarities on all 5000 instances for each value of $\tau$.
C.2.2 Simulation in Section 5
Table 1: The proportion of keep topic, ambiguous, and change of topic with varying $T_{test}=\{4,8,16,24,32,64,128,256,512\}$.

### v3

@@ 5 Explaining Topic Shifts
Definition 4 defines an ambiguous sequence as one where the highest-probability next-token predictions include tokens from both within and outside the input topic, reflecting natural ambiguity from overlapping topics. Take the second input sequence ${\bf X}=[{\bf e}_{1},{\bf e}_{4},{\bf e}_{1},{\bf e}_{4}]^{\top}$ in Figure 3 (right) as an example. $\widehat{\mathcal{G}}_{a}^{(4)}({\bf X})$ is $\{{\bf e}_{1}\}$, as depicted in $\mathcal{G}_{a}^{(4)}$ from Figure 3 (left) and $\widehat{\mathcal{G}}_{ab}^{(4)}({\bf X})$ is $\{{\bf e}_{1},{\bf e}_{4}\}$, as shown in $\mathcal{G}_{ab}^{(4)}$ from Figure 3 (left). $\widehat{\mathcal{G}}_{a}^{(4)}({\bf X})$ is a subset of $\widehat{\mathcal{G}}_{ab}^{(4)}({\bf X})$, although $\hat{y}_{{\bf w}_{ab}}\not\in\widehat{\mathcal{G}}_{a}^{(4)}({\bf X})$. We can argue that the next token predicted from an ambiguous sequence cannot be considered as a topic change, as it lacks the clear trigger phenomenon observed in human cognition. To address this, we propose a formal definition for a topic change:
@@
In Figure 3 (right), ${\bf W}_{ab}$ changes topic for the last input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{4},{\bf e}_{4},{\bf e}_{4}]^{\top}$, following the Definition 5. Building on the formal definitions of topic continuity, ambiguous sequences, and topic changes, we now present a necessary condition for a sequence to induce a topic change. This is achieved by introducing our final definition, grounded in the highest-priority SCC as determined by the order in the attention layer.
@@
Under the same settings and assumptions in Theorem 3, let ${\bf X}$ be a sequence that belongs to $\mathbb{T}_{a}$. If $\tilde{\bf W}_{ab}$ changes topic $\mathbb{T}_{a}$ for $\bf X$ then $\exists x_{j}\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$ such that $\forall x_{i}\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$, the number of times ${\bf x}_{j}$ appears in $\bf X$ is greater than the number of times ${\bf x}_{i}$ appears in $\bf X$.
Theorem 4 implies that, for a given sequence X from $\mathbb{T}_{a}$ and its corresponding TPG, a necessary condition for a topic change is the presence of a lower-priority token that appears more frequently than any of the higher-priority tokens. This can be intuitively understood through our analogy: if the oracle is following a conversation on Topic A but the conversation contains repeated components with lower importance in Topic A, its knowledge of Topic B may steer the response toward Topic B, thereby initiating a shift away from Topic A. A natural question arises: what do these findings imply in practice? Specifically, how does the probability of change of topic behave as the input sequence length or the topic ambiguity increases? The following theorem sheds light on these dynamics.
@@
    If $p>\max_{x\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})}\mathbb{P}\left({\bf x}_{i}^{\prime}={\bf x}\right)$,
    then $\lim_{T\to\infty}{\mathbb{P}(\tilde{\bf W}_{ab}\text{ changes topic }\mathbb{T}_{a}\text{ for }{\bf X}^{\prime})}=0$.
@@
    If $l$ increases then the probability that $\exists x_{j}^{\prime}\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$ such that $\forall x_{i}^{\prime}\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$, ${\bf x}_{j}^{\prime}$ outnumbers ${\bf x}_{i}^{\prime}$ in $\bf X^{\prime}$ does not increase.
@@
To illustrate Theorem 5 with simulations, we generate embeddings with $K=10$ and $d=16$. We approximate $\tilde{\mathbf{W}}_{a}$ and $\tilde{\mathbf{W}}_{ab}$ as the results obtained after $\tau=8000$ iterations of Algo-GD. We quantify the proportion of test sequences in which $\tilde{{\bf W}}_{ab}$ keeps $\mathbb{T}_{a}$ (keep topic), proportion of ambiguous sequences in $\mathbb{T}_{ab}$ (ambiguity) and proportion of times in which $\tilde{{\bf W}}_{ab}$ changes topic (change topic). First, we explore the effect of longer sequences by varying the length $T$ of the test sequences $\bf Z$. We increase $T$ from 4 to 512. Figure 5 illustrates how the proportion of change topic decreases as $T$ increases. Second, we investigate the effect of overlapping topics by generating TPGs with an increased number of edges $L$. Intuitively, a higher $L$ results in an increase $l$ and a greater overlap between TPGs of different topics. We vary $L$ from 4 to 18. Figure 5 demonstrates that as $L$ increases, ambiguity increases, while proportion of change topic doesn’t increase. These two findings contrast with expectations derived from human cognition but align with the result of Theorem 5. Lastly, among the 85000 test sequences generated for these experiments, 99.98% satisfy Theorem 4 (i.e., topic changes occur when a low-priority token appears more frequently than high-priority tokens). The remaining 0.02% mismatched cases were solely due to minor approximation discrepancies in the attention softmax. These results validate Theorem 4 (See simulation details in Appendix D).
(a) Input length.
(b) Topic ambiguity.
Figure 4: The proportion of topic continuity, ambiguity, and topic change as (a) input length and (b) topic ambiguity increase.
(a) Input length.
(b) Topic ambiguity.
Figure 5: Cosine similarity between continuations generated with single-topic and mixed-topic knowledge as (a) input length and (b) topic ambiguity vary.
6 Experiments in Frontier LLMs
To prove Theorem 5 we work within the simplified, single-layer self-attention model of Li et al., 2024b . Although this abstraction omits many hallmarks of contemporary LLMs (deep stacks of attention blocks, alternative cost functions, and other training heuristics), it offers a mathematically tractable setting that lets us derive interesting mathematical results. These results, in turn, can be used to understand how cutting-edge LLMs behave in terms of spontaneous topic changes. We empirically investigate such behavior on four frontier models: GPT-4o, Llama-3.3, Claude-3.7, and DeepSeek-V3.
Real dataset. We randomly select 100 arXiv papers published in March 2025 since the publicly disclosed knowledge cutoff dates for our study LLMs fall at the end of 2024 or earlier. This ensures that these models have not been trained on these data. We consider each paper as a different “topic”.
Experimental setup. For two distinct papers A and B, and an input prompt ($\bf X)$ from paper A, we consider a measure of topic continuity as the cosine similarity between the embeddings of the texts generated when the LLM has contextual knowledge solely from paper A ($\hat{y}_{\bf{W}_{a}}$) and when the LLM has contextual knowledge from both paper A and B ($\hat{y}_{\bf{W}_{ab}}$). We consider this cosine similarity as an empirical proxy for our definition of topic continuity (Definition 3): therefore the larger the similarity, the smaller the chance that the model has lead to a change of topic. This proxy suggests two testable consequences which become the empirical counterpart of our Theorem 5: Cosine similarity is expected to increase with the length of the input prompt, and is not expected to decrease when increasing ambiguity between paper A and paper B.
To more closely align with our theoretical framework, where a model gains knowledge of topic A and incrementally gains knowledge of topic B, we implement a Retrieval-Augmented Generation (RAG) approach, retrieving information exclusively from paper A or jointly from papers A and B (Von Oswald et al.,, 2023). Based on input prompt, we retrieve the top 3 most relevant excerpts from paper A or B to form the contextual knowledge set A or set B. The combined contextual knowledge set is simply the union of sets A and B. We add set A to the input prompt to obtain the generated text with sole knowledge of paper A ($\hat{y}_{\bf{W}_{a}}$), and we add the combined set to the input prompt to obtain the generated text with combined knowledge of paper A and B ($\hat{y}_{\bf{W}_{ab}}$). To closely follow our greedy decoding approach in our theoretical framework, we set the temperature parameter to 0 for all LLMs.
We designate each paper as paper A, and we randomly select 5 different papers from the remaining 99 papers as distinct paper B. For each input segment, we calculate the average cosine similarity between $\hat{y}_{\bf{W}a}$ and $\hat{y}_{\bf{W}{ab}}$ across these five pairs of paper A and paper B, using each LLM. The results for each LLM are averaged over the 100 papers (See more experimental details in Appendix E).
Experiment 1: Impact of the input length. We use the first $10,30,\ldots,150$ words from each paper A’s abstract as the input prompt. Figure 5 plots, for each LLM, the average cosine similarity as a function of the input length; the shaded bands indicate 95% confidence intervals. Across all models, similarity tends to increase with prompt length, aligning with the behavior predicted by Theorem 5.
Experiment 2: Impact of the topic ambiguity. We fix the input prompt length to the first $80$ words for each paper A’s abstract. We quantify paper ambiguity by the average similarity among each paper’s keywords: higher keyword similarity signifies lower paper ambiguity because topic is more specific. We partitioned the papers into six equal-width bins along this ambiguity spectrum. Figure 5 summarises the results: each boxplot shows the distribution of cosine similarities within an ambiguity bin (x-axis ordered from least to most ambiguous). Across all LLMs the median similarity does not seem to decrease, in agreement with the prediction of Theorem 5.
Taken together, the two experiments provide preliminary empirical support for Theorem 5, showing that its prediction, derived for a single-layer self-attention toy model, can be extended to today’s deep, multi-layer LLMs. Crucially, an important divergence between machine and human cognition persists in these frontier models: neither longer prompts nor greater topic ambiguity appreciably raise the likelihood of an spontaneous topic change.
7 Related Work
@@
Training and generalization of Transformer. (1) Properties of Softmax. The self-attention mechanism employs the softmax function to selectively emphasize different parts of the input. Gu et al., (2024), Goodfellow et al., (2016), and Deng et al., (2023) underscore the pivotal role of the softmax function in shaping attention distributions, influencing how models process and prioritize information within input sequences. Bombari and Mondelli, (2024) examined the word sensitivity of attention layers, revealing that softmax-based attention-layers are adept at capturing the significance of individual words. However, recent work has also pointed out limitations of the softmax function (Saratchandran et al.,, 2024; Deng et al.,, 2023). (2) Optimization in attention-based models. Additionally, recent researches interpret Transformer models as kernel machines, akin to support vector machines (SVMs), with self-attention layers performing maximum margin separation in the token space (Tarzanagh et al., 2023a, ; Tarzanagh et al., 2023b, ; Li et al., 2024b, ; Julistiono et al.,, 2024). (3) Chain-of-Thought (CoT) and In-Context Learning (ICL). Moreover, transformers exhibit remarkable abilities in generalization through ICL, where models effectively learn from contextual cues during inference (Brown et al.,, 2020; Xie et al.,, 2022; Olsson et al.,, 2022). CoT prompting (Wei et al.,, 2022; Zhou et al.,, 2023; Shao et al.,, 2023; Li et al., 2024a, ) enhances this by breaking down reasoning processes into intermediate steps, highlighting the emergent reasoning abilities of transformers. (4) Improvement efficiency of transformers. Recent advancements aim to improve the computational efficiency of transformers (Kitaev et al.,, 2020; Choromanski et al.,, 2021; Sukhbaatar et al.,, 2019; Wang et al.,, 2020), ensuring their viability for large-scale deployment while maintaining or enhancing their representational capabilities.
@@
Self-Attention and topic dynamics. Advancements in self-attention research have deepened our understanding of how transformers handle evolving semantic contexts. Prior work has explored diverse aspects of topic modeling, such as dynamic topic structures (Miyamoto et al.,, 2023), hierarchical relationships (Lin et al.,, 2024), topic-aware attention mechanisms (Panwar et al.,, 2021), and the mechanistic underpinnings of topic representation (Li et al.,, 2023). While these studies provide insights into managing static and hierarchical topic structures, our work focuses on the topic changes with the given input sequences from a specific topic.
@@ 8 Discussion
Our theoretical analysis on self-attention models and empirical investigations on modern LLMs reveal fundamental clues regarding the distinctions between model-based spontaneous topic changes and human spontaneous thought, a phenomenon that is critical for comparing conversational dynamics across humans and AI. In an era of growing concern about AI’s cognitive resemblance to humans, our framework provides preliminary results differentiating these phenomena, thereby opening pathways for future interdisciplinary research at the interface of artificial and human cognition.
Limitations. Our theoretical framework builds on the same simplified single-layer self-attention model with a log-loss objective from Li et al., 2024b and defining topics as TPGs. These abstractions do not fully capture the complexities of contemporary LLMs, including deep attention architectures, alternative loss functions, and diverse training objectives. Despite loosening these assumptions, our experiments suggest that the essence of our core theoretical conclusions holds across modern LLMs within our framework of study. Future work will investigate how broadly these theoretical insights generalize to complex LLM architectures.
@@ Appendix A Technical Proofs
A.1 Proof of Lemma 2
@@ Appendix B Detailed Explanation of Assumption 4
Figure 6: Illustration of Assumption 4. Here are two datasets related to the TPGs, $\mathcal{G}_{a}^{(4)}$ and $\mathcal{G}_{b}^{(4)}$, from Figure 3 (left). From the directed arrows in $\mathcal{G}_{a}^{(4)}$, we can generate a dataset with the last token $\bf{e}_{4}$ for $\mathbb{T}_{a}$, which can reconstruct back the $\mathcal{G}_{a}^{(4)}$. A similar process applies for the $\mathcal{G}_{b}^{(4)}$.
Appendix C Explanation of TPGs Combination
Figure 3 presents $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$ as the TPGs corresponding to the last input token ${\bf e}_{4}$ for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, respectively. In $\mathcal{G}^{(4)}_{a}$, the token priority is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{2}>{\bf e}_{4}$. In $\mathcal{G}^{(4)}_{ab}$ for the mixed-topics, the priority order is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{2}={\bf e}_{4}$. The strict equality ${\bf e}_{1}={\bf e}_{2}$ from $\mathcal{G}^{(4)}_{a}$ is maintained in $\mathcal{G}^{(4)}_{ab}$, but the strict inequality ${\bf e}_{2}>{\bf e}_{4}$ is replaced with ${\bf e}_{2}={\bf e}_{4}$ in mixed-topics, forming the new SCC, $\{{\bf e}_{1},{\bf e}_{2},{\bf e}_{4}\}$, in $\mathcal{G}^{(4)}_{ab}$.
Appendix D Detailed Simulation Studies with Single-layer Self-attention
D.1 Simulation Process
@@
D.2 Additional Experiments to Support Theorem 3
The attention priority similarity quantifies how well the weights ${\bf W}^{\prime}$ preserve the attention priority of the weights ${\bf W}$. A value of 1 indicates that the priority is fully preserved. Using this metric, we conduct experiments, with results in Figure 7. We generate embeddings with $K=10$ and $d=16$, and randomly construct TPGs for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$. Using these TPGs, we randomly generate $\text{DSET}_{a}$ and $\text{DSET}_{b}$. We compute $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$, $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ and $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ using the same procedure as Li et al., 2024b . We generate test sequences ${\bf Z}$ within $\mathbb{T}_{a}$, and we calculate the attention priority similarity of $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ relative to both $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ and $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$. We repeat this process for multiple TPGs and input sequences (simulation details in Appendix D). Figure 7 clearly demonstrates that the similarity converges to 1 after $\tau=8000$ iterations when evaluated relative to $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (blue line), but fails to converge relative to $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ (orange line). These observations align with the results of Theorem 3.
Figure 7: Convergence of attention priority similarity for $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ relative to $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (blue) and $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ (orange).
D.3 Simulation in Section 5
Table 1: Proportion of keep topic, ambiguous, and change of topic with varying $T_{test}=\{4,8,16,24,32,64,128,256,512\}$.